### v1

# On an $n$-ary generalization of the Lie representation

Abstract.
We continue our study, initiated in our prior work with Richard Stanley, of the representation of the symmetric group on the multilinear component of an $n$-ary generalization of the free Lie algebra known as the free Fillipov $n$-algebra with $k$ brackets. Our ultimate aim is to determine the multiplicities of the irreducible representations in this representation. This had been done for the ordinary Lie representation ($n=2$ case) by Kraskiewicz and Weyman. The $k=2$ case was handled in our prior work, where the representation was shown to be isomorphic to $S^{2^{n-1}1}$. In this paper, for general $n$ and $k$, we obtain decomposition results that enable us to prove that in the $k=3$ case, the representation is isomorphic to $S^{3^{n-1}1}\oplus S^{3^{n-2}21^{2}}$.
${}^{2}$upported in part by NSF grants DMS 1502606 and DMS 2207337
1. Introduction
There are different generalizations of the notion of Lie algebra that involve replacing the binary Lie bracket with an $n$-ary bracket. One was introduced by Filippov in [Fi] and another was introduced by Hanlon and Wachs in [HW]. For both generalizations, the representation of the symmetric group on the multilinear component of the free $n$-ary Lie algebra generalizes a well studied representation known as the Lie representation; see e.g. [Ba, KW, Re, St, Wa] for work on the Lie representation. The representation for the Hanlon-Wachs generalization was first studied in [HW]. More recently, a study of the representation for the Filippov generalization was initiated by Friedmann, Hanlon, Stanley and Wachs in [FHSW1]. In this paper, we continue this study.
Throughout this paper, all vector spaces are taken over a field of characteristic 00, say ${\mathbb{C}}$. Fix $n\geq 2$. A Filippov algebra is a vector space ${\cal L}$ equipped with an $n$-linear map (called a bracket)
[,,,]:×$[\cdot,\cdot,\hskip 19.91684pt,\cdot]:\times^{n}{\cal L}\rightarrow{\cal L}~{}$
that satisfies the following antisymmetry relation for all $\sigma$ in the symmetric group $S_{n}$:
(1.1) $[x_{1},\dots,x_{n}]=\operatorname{sgn}(\sigma)[x_{\sigma(1)},\dots,x_{\sigma(n)}]$
and the following generalization of the Jacobi identity:
(1.2) $[[x_{1},x_{2},\ldots,x_{n}],x_{n+1},\ldots,x_{2n-1}]$
$=\sum_{i=1}^{n}[x_{1},x_{2},\ldots,x_{i-1},[x_{i},x_{n+1},\ldots,x_{2n-1}],x_{i+1},\ldots,x_{n}],$
for $x_{i}\in{\cal L}$. When we need to specify $n$, we say Filippov $n$-algebra. Clearly Filippov $2$-algebras are the same as ordinary Lie algebras. In [FHSW1], Filippov $n$-algebras are referred to as Lie algebras of the nth kind (LAnKe’s for short, see [Fr]).
As defined in [FHSW1], the free Filippov $n$-algebra (or free LAnKe) on a set $X=\{x_{1},x_{2},\dots,x_{m}\}$ is a vector space generated by all possible $n$-ary bracketings of elements of $X$, where the only possible relations existing among these bracketings are consequences of $n$-linearity of the bracketing, the antisymmetry of the bracketing (1.1) and the generalized Jacobi identity (1.2).
The multilinear component of the free Filippov $n$-algebra on $X$ is the subspace spanned by $n$-bracketed permutations of $X$. Every such $n$-bracketed permutation on $X$ has the same number of brackets, which depends on $m$, the number of generators. Indeed, if the number of brackets is $k$ then $m=(n-1)k+1$.
The object we studied in [FHSW1] and continue to study in this paper is the representation of the symmetric group $S_{(n-1)k+1}$ on the multilinear component of the free Filippov $n$-algebra on $(n-1)k+1$ generators. We denote this representation by $\rho_{n,k}$ and note that $\rho_{2,k}$ is the classical Lie representation $\mathrm{Lie}_{k+1}$. It follows from the antisymmetry of the bracket that when $k=1$, $\rho_{n,k}$ is the sign representation $\operatorname{sgn}_{n}$ which is the same as the Specht module $S^{1^{n}}$. A result of Kraskiewicz and Weyman [KW] on $\mathrm{Lie}_{m}$ gives the decomposition of $\rho_{n,k}$ into irreducibles when $n=2$. Indeed, the multiplicity of the irreducible representation indexed by $\lambda$ in $\rho_{2,k}$ is given by the number $\tilde{f}^{\lambda}$ of standard Young tableaux of shape $\lambda$ and major index congruent to $11\bmod k+$ 1.
Table 1 below summarizes the decompositions of $\rho_{n,k}$ into irreducibles that were previously established or will be established in this paper. The partitions $\lambda$ in the table stand for Specht modules $S^{\lambda}$.
$n\backslash k$ $1$ $2$ $3$ $4$ $k$
$2$ $1^{2}$ $21$ $31\oplus 21^{2}$ $41\oplus 32\oplus 31^{2}\oplus 2^{2}1\oplus 21^{3}$ $\oplus_{\lambda\vdash k+1}\tilde{f}^{\lambda}S^{\lambda}$
$3$ $1^{3}$ $12^{2}$ 1 $3^{2}1\oplus 321^{2}$ $4^{2}1\oplus 432\oplus 431^{2}\oplus 42^{2}1\oplus 421^{3}\oplus\gamma_{3,4}$111where $\gamma_{3,4}$ is an $S_{9}$-module of dimension $204$ whose irreducibles $S^{\lambda}$ have $\lambda_{1}\leq 3$. $\rho_{3,k}$
$4$ $1^{4}$ $12^{3}$ 1 $3^{3}1\oplus 3^{2}21^{2}$ $\rho_{4,4}$ $\rho_{4,k}$
$n$ $1^{n}$ $12^{n-1}$ 1 $3^{n-1}1\oplus 3^{n-2}21^{2}$ $\rho_{n,4}$ $\rho_{n,k}$
Table 1. What we now know about $\rho_{n,k}$.
In [FHSW1], the $k=2$ column of the table was established with the following result.
Theorem 1.1 ([FHSW1]).
For all $n\geq 2$, the $S_{2n-1}$-module $\rho_{n,2}$ is isomorphic to the Specht module $S^{2^{n-1}1}$, whose dimension is the $n^{th}$ Catalan number $\frac{1}{n+1}{2n\choose n}$.
In this paper we establish the $k=3$ column of the table with the following result.
Theorem 1.2.
For all $n\geq 2$, as $S_{3n-2}$-modules,
$\rho_{n,3}\cong S^{3^{n-1}1}\oplus S^{3^{n-2}21^{2}}.$
Consequently, by the hook-length formula, $\dim\rho_{n,3}=\displaystyle\frac{4}{\prod_{i=1}^{3}(n+i)}{3n\choose n,n,n}$.
We see that for $k\in\{1,2,3\}$ and $n\geq 2$, increasing $n$ by 1 adds a row of length $k$ to the top of each Young diagram indexing the Specht module in the decomposition of $\rho_{n-1,k}$. More precisely, for $k\in\{1,2,3\}$, we have
(1.3) $\rho_{n,k}\cong\beta_{n,k}\iff n\geq k,$
where $\beta_{n,k}$ is the $S_{(n-1)k+1}$-module whose decomposition into irreducibles is obtained by adding a row of length $k$ to the top of each Young diagram in the decomposition of $\rho_{n-1,k}$. Here we set $\beta_{1,k}=\rho_{1,k}=S^{1}$. We leave open the question of whether (1.3) is true for all $k$. The Kraskiewicz-Weyman decomposition of $\rho_{2,k}=\mathrm{Lie}_{k+1}$ and the $(3,4)$ entry of Table 1, as well as columns $k=1,2,3$, provide evidence of an affirmative answer. We provide additional evidence in this paper with the following general result. See [FHSW1, Sec. 4] for further evidence.
Theorem 1.3.
Let $n,k\geq 1$. Then as $S_{(n-1)k+1}$-modules,
(1.4) $\rho_{n,k}\cong\beta_{n,k}\oplus\gamma_{n,k},$
for some $S_{(n-1)k+1}$-module $\gamma_{n,k}$ all of whose irreducibles have Young diagrams with at most $k-1$ columns.
Remark 1.4.
Theorems 1.2 and 1.3 were announced in our extended abstract [FHSW0] and in our paper [FHSW1], with a reference to our current paper (which was in preparation at the time of publication of [FHSW1]). Maliakas and Stergiopoulou have recently posted a preprint (arXiv:2401.09405) giving a proof of Theorem 1.2 that appears to be very different from ours.
This paper is organized as follows. Some preliminary definitions and results are given in Section 2. In Section 3, we prove Theorem 1.3, which is then used to obtain the decomposition for $\rho_{3,4}$ given in Table 1. We also present in Section 3, a result on the submodule of $\rho_{n,k}$ spanned by non-combs, which is used in the proof of Theorem 1.2 given in Section 4.
2. Preliminaries
2.1. Specht modules
The Specht modules $S^{\lambda}$, where $\lambda$ is a partition of $n$, give a complete set of irreducible representations of the symmetric group $\mathfrak{S}_{n}$ over a field of characteristic 00, say $\mathbb{C}$. They can be constructed as subspaces of the regular representation $\mathbb{C}\mathfrak{S}_{n}$ or as presentations given in terms of generators and relations, known as Garnir relations. We will need the latter type of construction in this paper.
Let $\lambda=(\lambda_{1}\geq\dots\geq\lambda_{l})$ be a partition of $n$. A Young tableau of shape $\lambda$ is a filling of the Young diagram of shape $\lambda$ with distinct entries from the set $[n]:=\{1,2,\dots,n\}$. Let $\mathcal{T}_{\lambda}$ be the set of Young tableaux of shape $\lambda$.
To construct the Specht module as a presentation, one can use column tabloids and Garnir relations. Let $M^{\lambda}$ be the vector space (over ${\mathbb{C}}$) generated by $\mathcal{T}_{\lambda}$ subject only to column relations, which are of the form $t+s$, where $s\in\mathcal{T}_{\lambda}$ is obtained from $t\in\mathcal{T}_{\lambda}$ by switching two entries in the same column. Given $t\in\mathcal{T}_{\lambda}$, let $\bar{t}$ denote the coset of $t$ in $M^{\lambda}$. These cosets, which are called column tabloids, generate $M^{\lambda}$. A Young tableau is column strict if the entries of each of its columns increase from top to bottom. Clearly, $\{\bar{t}:t\mbox{ is a column strict Young tableau of shape }\lambda\}$ is a basis for $M^{\lambda}$.
We will use a presentation of $S^{\lambda}$ discussed in Fulton [Fu, page 102 (after Ex. 15)]. The generators are the column tabloids $\bar{t}$, where $t\in\mathcal{T}_{\lambda}$. The Garnir relations are of the form [^2] In our previous papers [FHSW1, FHW], we denote $g^{t}_{c}$ by $g^{t}_{c,1}$.
$g^{t}_{c}:=\bar{t}-\sum_{s}\bar{s},$
where $c\in[\lambda_{1}-1]$, $t\in\mathcal{T}_{\lambda}$, and the sum is over all $s\in\mathcal{T}_{\lambda}$ obtained from $t\in\mathcal{T}_{\lambda}$ by exchanging any entry of column $c$ with the top entry of the next column.
Example: For
[…]
we have
[…]
Permutations $\sigma$ in $\mathfrak{S}_{n}$ act on a tableau $t$ by replacing each entry of $t$ with its image under $\sigma$. Note that this action respects the column relations and the Garnir relations.
Proposition 2.1 ([Fu]).
For all $\lambda\vdash n$, the Specht module $S^{\lambda}$ is generated by the column tabloids subject only to the Garnir relations $g^{t}_{c}$, where $t\in\mathcal{T}_{\lambda}$ and $c\in[\lambda_{1}-1]$.
2.2. Filippov Algebras
In Appendix 1 of [DI], a proof that the generalized Jacobi relations (1.2) are equivalent to the relations
(2.1) $[[x_{1},x_{2},\ldots,x_{n}],y_{1},\ldots,y_{n-1}]$
$=$ $\sum_{i=1}^{n}[[x_{1},x_{2},\ldots,x_{i-1},y_{1},x_{i+1},\ldots,x_{n}],x_{i},y_{2},\ldots,y_{n-1}]$
is given. This gives an alternative presentation of $\rho_{n,k}$ for all $n,k$.
An $n$-bracketed word on a set $A$ is called an $n$-comb (or just comb) if it has the form
$[[\cdots[[a_{0},a_{1,1},\dots,a_{1,n-1}],a_{2,1},\dots,a_{2,n-1}],\dots],a_{k,1},\dots,a_{k,n-1}],$
where $a_{0},a_{i,j}\in A$.
Proposition 2.2.
Let $Lie_{n,k}(X)$ be the component of the free Filippov $n$-algebra on a set $X$ spanned by $n$-bracketed words with $k$ brackets. Then the combs in $Lie_{n,k}(X)$ span $Lie_{n,k}(X)$.
Proof.
Let $w$ be a bracketed word in the free Filippov $n$-algebra. We will prove by induction on the number $b(w)$ of brackets of $w$ that $w$ is a linear combination of combs. If $b(w)=0,1$ then $w$ is clearly a comb. Now let $b(w)\geq 2$ and $w=[w_{1},w_{2},\dots,w_{n}]$. Define degree $d(w):=\max_{i\in[n]}b(w_{i})$.
We claim that if $w$ has the highest possible degree $b(w)-1$ then $w$ is a linear combination of combs. Indeed, in this case all but one $w_{i}$ is a single letter. Using antisymmetry of the bracket, we can assume without loss of generality that $b(w_{1})=b(w)-1$ and $b(w_{i})=0$ for $i=2,\dots,n$. By induction $w_{1}$ is a linear combination of combs, which implies by linearity of the bracket that $w$ is a linear combination of combs.
Now let $d(w)<b(w)-1$. We only need to show that $w$ is equal to a linear combination of bracketed words of higher degree. Clearly there are at least two $w_{i}$ for which $b(w_{i})>0$. . By antisymmetry of the bracket, without loss of generality, we can assume that $0<b(w_{1})\leq b(w_{2})$ and that $d(w)=b(w_{2})$. Let $w_{1}=[u_{1},\dots,u_{n}]$. Using the alternative generalized Jacobi relation given in (2.1), we have
$w=\sum_{i=1}^{n}[[u_{1},u_{2},\dots,u_{i-1},w_{2},u_{i+1},\dots,u_{n}],u_{i},w_{3},w_{4},\dots,w_{n}].$
Note that the degree of each term is greater than or equal to $b(w_{2})+1=d(w)+1$. Thus $w$ is a sum of bracketed words of degree greater than $d(w)$. ∎
Let $V_{n,k}$ be the multilinear component of the vector space generated by $n$-combs on $[k(n-1)+1]$ subject only to the anticommuting relations, but not the generalized Jacobi relations. Clearly $V_{n,k}$ is invariant under the action of $S_{k(n-1)+1}$ on the multilinear component of the full vector space of bracketed permutations on $[k(n-1)+1]$ subject only to the anticommuting relations. By Proposition 2.2, $\rho_{n,k}$ is a quotient (or submodule) of $V_{n,k}$. Note that as $S_{k(n-1)+1}$-modules,
(2.2) $V_{n,k}\cong M^{n^{1}(n-1)^{k-1}}\cong(\operatorname{sgn}_{n}\otimes\operatorname{sgn}_{n-1}^{k-1})\uparrow_{S_{n}\times S_{n-1}^{k-1}}^{S_{k(n-1)+1}}.$
As a consequence, by Pieri’s rule, we have the following useful result.
Proposition 2.3.
Let $\lambda=(\lambda_{1}\geq\dots\geq\lambda_{\ell})$ be a partition of $k(n-1)+1$. If $S^{\lambda}$ is in $\rho_{n,k}$ then $\lambda_{1}\leq k$.
Throughout this paper, given a partition $\lambda$, we will denote its largest part by $\lambda_{1}$, (which is the same as the number of columns in the Young diagram associated with $\lambda$).
3. A general decomposition
In this section we prove Theorem 1.3 and then use it to prove that $\rho_{3,4}$ has the decomposition given Table 1. We also present a result relating $\rho_{n,k}$ to the submodule of $\rho_{n,k}$ spanned by non-combs, which will be used in the proof of Theorem 1.2 given in Section 4.
Let $\hat{\rho}_{n,k}$ be the subspace of the free Filippov $n$-algebra on ${[k(n-2)+2]}$ spanned by bracketed words of type $1^{k(n-2)+1}k$, where the type of a bracketed word is the weak composition $(\mu_{1},\mu_{2},\dots,\mu_{k(n-2)+2})$ with $\mu_{i}$ equaling the number of $i$’s in the bracketed word. This means that the letters $11,2,\dots,k(n-2)+$ 1 each occur once in the bracketed words generating $\hat{\rho}_{n,k}$ and the letter $b:=k(n-2)+2$ occurs $k$ times. The symmetric group $S_{k(n-2)+1}$ acts on $\hat{\rho}_{n,k}$ by replacing each letter $11,\dots,k(n-2)+$ 1 in a bracketed word with its image under a permutation in $S_{k(n-2)+1}$, leaving $b$ fixed.
We also need an extended definition of Specht modules that allows for repeated entries. We define a tableau to be a filling of a shape $\lambda$ with positive integer entries. The type of a tableau $t$ is the weak composition whose $i$th entry is the number of $i$’s appearing in $t$. For each partition $\lambda$ of $k(n-1)+1$, let $\hat{\mathcal{T}}_{\lambda}$ be the set of tableaux of shape $\lambda$ and type $1^{k(n-2)+1}k$. The column relations, column tabloids, and Garnir relations on $\hat{\mathcal{T}}_{\lambda}$ are the same as on $\mathcal{T}_{\lambda}$ given in Section 2. Now let $\hat{S}^{\lambda}$ be the vector space generated by column tabloids $\bar{t}$, where $t\in\hat{\mathcal{T}}_{\lambda}$, subject to the Garnir relations $g^{t}_{c}$. The symmetric group $S_{k(n-2)+1}$ acts on $\bar{t}$ by replacing each entry $11,\dots,k(n-2)+$ 1 of $t$ with its image under a permutation in $S_{k(n-2)+1}$, leaving $b$ fixed.
In [DI, Proposition 43], the idea of obtaining Filippov $(n-1)$-algebras from Filippov $n$-algebras is discussed. We use this idea in the proof of the following result.
Lemma 3.1.
Let $n\geq 2$ and $k\geq 1$. Then
(3.1) $\hat{\rho}_{n,k}\cong_{S_{k(n-2)+1}}\rho_{n-1,k},$
and for all $\lambda\vdash k(n-1)+1$ with $\lambda_{1}=k$,
(3.2) $\hat{S}^{\lambda}\cong_{S_{k(n-2)+1}}S^{\lambda^{-}},$
where if $\lambda$ is the partition $(\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{\ell})$ then $\lambda^{-}$ is the partition $(\lambda_{2}\geq\dots\geq\lambda_{\ell})$.
Proof of (3.1).
The $n$-ary bracketed words of type $1^{k(n-2)+1}k$ have $k$ occurences of the letter $b:=k(n-1)+2$. Since there are $k$ brackets, by antisymmetry of the brackets, each bracket of a nonvanishing $n$-ary bracketed word contains exactly one $b$ and the $b$ can be placed at the right end of each bracket. For example,
$[[[3,1,b],[5,2,b],b],[6,4,b],b]$
is such a $3$-bracketed word of type $51^{5(1)+1}$ 5. These bracketed words, which we refer to as $b$-bracketings, generate $\hat{\rho}_{n,k}$. The antisymmetry relations reduce to
$[x_{1},\dots,x_{n-1},b]=\operatorname{sgn}(\sigma)[x_{\sigma(1)},\dots,x_{\sigma(n-1)},b],$
where $\sigma\in S_{n-1}$, for the $b$-bracketings, and the generalized Jacobi relations reduce to
$[[x_{1},x_{2},\ldots,x_{n-1},b],x_{n+1},\ldots,x_{2n-2},b]$
$=\sum_{i=1}^{n-1}[x_{1},x_{2},\ldots,x_{i-1},[x_{i},x_{n+1},\ldots,x_{2n-2},b],x_{i+1},\ldots,x_{n-1},b]$
for the $b$-bracketings, since the $i=n$ term of the summation in (1.2) is
$[x_{1},\dots,x_{n-1},[b,x_{n+1},\dots,x_{2n-2},b]],$
which by antisymmetry is 00. These relations imply all relations for $\hat{\rho}_{n,k}$. Indeed, one can check that any other relation on these generators that comes from (1.2) is trivial. For instance, by antisymmetry of the bracket, (1.2) with $x_{n+1}=x_{2n-1}=b$ yields the relation $0=0$ and (1.2) with $x_{1}=x_{n}=b$ yields the relation $0=[[x_{n+1},\dots,x_{2n-1},b],x_{2},\dots,x_{n-1},b]-[[x_{n+1},\dots,x_{2n-1},b],x_{2},\dots,x_{n-1},b]$.
Now let $\varphi:\hat{\rho}_{n,k}\to\rho_{n-1,k}$ be the linear map defined by letting $\varphi(w)$ be the $(n-1)$-ary bracketed word obtained from the $b$-bracketing $w$ by removing all the $b$’s. For example,
$\varphi([[[3,1,b],[5,2,b],b],[6,4,b],b])=[[[3,1],[5,2]],[6,4]].$
We can see that $\varphi$ respects the antisymmetry relations and the generalized Jacobi relations and is therefore a well defined linear map. Indeed, $\varphi$ takes
$[x_{1},\dots,x_{n-1},b]-\operatorname{sgn}(\sigma)[x_{\sigma(1)},\dots,x_{\sigma(n-1)},b]$
to
$[x_{1},\dots,x_{n-1}]-\operatorname{sgn}(\sigma)[x_{\sigma(1)},\dots,x_{\sigma(n-1)}]$
for all $\sigma\in S_{n-1}$ and $\varphi$ takes
$[[x_{1},x_{2},\ldots,x_{n-1},b],x_{n+1},\ldots,x_{2n-2},b]$
$-\sum_{i=1}^{n-1}[x_{1},x_{2},\ldots,x_{i-1},[x_{i},x_{n+1},\ldots,x_{2n-2},b],x_{i+1},\ldots,x_{n-1},b]$
to
$[[x_{1},x_{2},\ldots,x_{n-1}],x_{n+1},\ldots,x_{2n-2}]$
$-\sum_{i=1}^{n-1}[x_{1},x_{2},\ldots,x_{i-1},[x_{i},x_{n+1},\ldots,x_{2n-2}],x_{i+1},\ldots,x_{n-1}].$
Note that $\varphi$ can easily be inverted. Hence $\varphi$ is a vector space isomorphism. It is clear that $\varphi$ respects the action of $S_{k(n-2)+1}$; so it is an $S_{k(n-2)+1}$-module isomorphism. ∎
Proof of (3.2).
The proof is similar to the proof of (3.1). The column tabloids $\bar{t}$ of shape $\lambda$ and type $1^{k(n-2)+1}k$ have $k$ occurences of $b:=k(n-2)+2$. By antisymmetry of the columns, since $\lambda$ has $k$ columns, each column of $t$ must have exactly one $b$ and the $b$ can be placed at the bottom of the column. We call such a $t$ a $b$-tableaux. Let $\mathcal{T}^{b}_{\lambda}$ be the subset of $\hat{\mathcal{T}}_{\lambda}$ consisting of $b$-tableaux of shape $\lambda$. The column relations on $b$-tableaux are of the form $t+s$, where $s$ is obtained from $t$ by exchanging two non-$b$ entries in the same column. The column tabloids $\bar{t}$, where $t\in\mathcal{T}^{b}_{\lambda}$, generate $\hat{S}^{\lambda}$. The Garnir relations for these generators reduce to $g^{t}_{c}=\bar{t}-\sum_{s}\bar{s}$, where $t\in\mathcal{T}^{b}_{\lambda}$ and the sum is over all $s$ obtained from $t$ by exchanging any element of column $c$ other than the bottom element $b$ with the top element of the next column (provided the next column has length greater than 1).
Given a $b$-tableaux $t$ of shape $\lambda$, let $t^{-}$ be the tableaux obtained from $t$ by removing the $b$ from the bottom of each column. Clearly $t^{-}$ is a tableaux of shape $\lambda^{-}$. Now let $\psi:\hat{S}^{\lambda}\to S^{\lambda^{-}}$ be the linear map defined on the generating set $\{\bar{t}\mid t\in\mathcal{T}^{b}_{\lambda}\}$ by letting $\psi(\bar{t})=\overline{t^{-}}$. To see that this is a well-defined $S_{k(n-2)+1}$-module homomorphism, one can easily verify that the column relations map to column relations, and the Garnir relation $g^{t}_{c}$ maps to the Garnir relation $g^{t^{-}}_{c}$, for all $t\in\mathcal{T}^{b}_{\lambda}$. The inverse map is easily obtained by adding a $b$ to the bottom of each column of a tableau of shape $\lambda^{-}$. Again since relations map to relations under this map, the inverse is well-defined. Hence $\psi$ is an isomorphism. Alternatively, one can see that $\psi$ is an isomorphism, by noting that $\psi$ maps basis elements of $\hat{S}^{\lambda}$ to basis elements of $S^{\lambda^{-}}$. Indeed, the column tabloids $\bar{t}$, where $t$ is a semistandard $b$-tableau of shape $\lambda$, form a basis for $\hat{S}^{\lambda}$ and the column tabloids $\bar{t}$, where $t$ is a standard tableaux of shape $\lambda^{-}$ form a basis for $S^{\lambda^{-}}$
Lemma 3.2.
As $S_{k(n-2)+1}$-modules
$\hat{\rho}_{n,k}\cong\bigoplus_{\begin{subarray}{c}\lambda\vdash k(n-1)+1\\ \lambda_{1}=k\end{subarray}}c_{\lambda}\hat{S}^{\lambda},$
where $c_{\lambda}$ is the multiplicity of the irreducible $S_{k(n-1)+1}$-module $S^{\lambda}$ in $\rho_{n,k}$.
Proof.
Let $B:=\{k(n-2)+2,k(n-2)+3,\dots,k(n-1)+1\}$ and let $\alpha:=\sum_{\sigma\in S_{B}}\sigma$. One can see that
$\alpha(\rho_{n,k}\downarrow^{S_{k(n-1)+1}}_{S_{k(n-2)+1}\times S_{B}})\cong\hat{\rho}_{n,k}\otimes 1_{S_{B}}$
and
$\alpha(S^{\lambda}\downarrow^{S_{k(n-1)+1}}_{S_{k(n-2)+1}\times S_{B}})\cong\hat{S}^{\lambda}\otimes 1_{S_{B}},$
where $1_{S_{B}}$ denotes the trivial representation of $S_{B}$. Now take the restriction of both sides of the decomposition
$\rho_{n,k}\cong\bigoplus_{\lambda\vdash k(n-1)+1}c_{\lambda}S^{\lambda},$
from $S_{k(n-1)+1}$ to $S_{k(n-2)+1}\times S_{B}$ and then take the image under the action of $\alpha$ to get the $S_{k(n-2)+1}$-module isomorphism
$\hat{\rho}_{n,k}\cong\bigoplus_{\lambda\vdash k(n-1)+1}c_{\lambda}\hat{S}^{\lambda}.$
By Proposition 2.3, $c_{\lambda}=0$ if $\lambda_{1}>k$. On the other hand $\hat{S}^{\lambda}=0$ if $\lambda_{1}<k$. Indeed, if $t\in\hat{\mathcal{T}}_{\lambda}$ then entry $b:=k(n-2)+2$ is repeated $k$ times. If $\lambda_{1}<k$ then entry $b$ will appear more than once in some column of $t$. By antisymmetry of the columns, the column tabloid $\bar{t}$ is equal to 00. Thus we can limit the summation to $\lambda$ with $\lambda_{1}=k$. ∎
We can now easily prove Theorem 1.3, which we restate here. First recall that for $n\geq 2$ and $k\geq 1$, we defined $\beta_{n,k}$ to be the $S_{k(n-1)+1}$-module whose decomposition into irreducibles is obtained from the $S_{k(n-2)+1}$-module $\rho_{n-1,k}$ by adding a row of length $k$ to the top of each Young diagram in the decomposition of $\rho_{n-1,k}$. Also recall that for $k\geq 1$, we defined $\rho_{1,k}$ to be $S^{1}$. For example,
$\rho_{1,3}$ $=S^{1}\mbox{ and }\beta_{2,3}=S^{21}$
$\rho_{2,3}$ $=S^{31}\oplus S^{21^{2}}\mbox{ and }\beta_{2,4}=S^{3^{2}1}\oplus S^{321^{2}}.$
Theorem 3.3 (Theorem 1.3).
Let $n,k\geq 1$. Then as $S_{(n-1)k+1}$-modules,
(3.3) $\rho_{n,k}\cong\beta_{n,k}\oplus\gamma_{n,k},$
for some $S_{(n-1)k+1}$-module $\gamma_{n,k}$ all of whose irreducibles have Young diagrams with at most $k-1$ columns.
Proof.
Lemmas 3.2 and 3.1 together imply that as $S_{k(n-2)+1}$-modules,
$\rho_{n-1,k}\cong\bigoplus_{\begin{subarray}{c}\lambda\vdash k(n-1)+1\\ \lambda_{1}=k\end{subarray}}c_{\lambda}S^{\lambda^{-}},$
where $c_{\lambda}$ is the multiplicity of $S^{\lambda}$ in $\rho_{n,k}$. If we add a part of size $k$ to $\lambda^{-}$, we get $\lambda$. Thus
$\beta_{n,k}\cong\bigoplus_{\begin{subarray}{c}\lambda\vdash k(n-1)+1\\ \lambda_{1}=k\end{subarray}}c_{\lambda}S^{\lambda}.$
The result now follows from Proposition 2.3. ∎
In [FHSW1, Section 4], we speculate that $\gamma_{n,k}=0$ if and only if $n\geq k$. The following result provides some evidence for this.
Corollary 3.4.
The $S_{9}$-module $\rho_{3,4}$ decomposes as
$\rho_{3,4}\cong S^{4^{2}1}\oplus S^{432}\oplus S^{431^{2}}\oplus S^{42^{2}1}\oplus S^{421^{3}}\oplus\gamma_{3,4},$
where $\gamma_{3,4}$ is an $S_{9}$-module of dimension $204$ that decomposes into irreducibles whose Young diagram has fewer than $4$ columns.
Proof.
From the Kraskiewicz-Weyman result [KW] discussed in the introduction (see Table 1),
$\rho_{2,4}\cong S^{4^{1}1}\oplus S^{32}\oplus S^{31^{2}}\oplus S^{2^{2}1}\oplus S^{21^{3}}.$
It follows that
$\beta_{3,4}\cong S^{4^{2}1}\oplus S^{432}\oplus S^{431^{2}}\oplus S^{42^{2}1}\oplus S^{421^{3}}.$
Using a computer program written in C++, we found that $\dim\rho_{3,4}=1077$. It follows from the hook length formula that $\dim\beta_{3,4}=873$. Hence the result follows from Theorem 3.3. ∎
Remark 3.5.
In [FHSW1, Section 4], we speculate that $\gamma_{3,4}\cong S^{3^{2}1^{3}}\oplus S^{32^{3}}$ and we provide some evidence for this.
Theorem 3.6.
For all $k,n\geq 1$, let $\tilde{\rho}_{n,k}$ be the submodule of $\rho_{n,k}$ spanned by non-combs. Then
  1. (1)
    the multiplicity of $S^{k^{n-1}1}$ in $\rho_{n,k}$ is 1
  2. (2)
    $\rho_{n,k}\cong S^{k^{n-1}1}\oplus\tilde{\rho}_{n,k}$
  3. (3)
    $\tilde{\rho}_{n,k}=0$ if and only if $k=1,2$.
Proof.
(1) follows immediately from Theorem 3.3 and induction on $n$. It is clearly true when $n=1$.
We prove (2) by constructing an $S_{k(n-1)+1}$-module isomorphism
$\psi:S^{k^{n-1}1}\to\rho_{n,k}/\tilde{\rho}_{n,k}.$
Let $t$ be the tableaux whose first column from top down is $a_{0},a_{1,1},\dots,a_{1,n-1}$ and whose $i$th column, where $2\leq i\leq k$, from top down is $a_{i,1},a_{i,2},\dots,a_{i,n-1}$. We define $\psi$ on generators of $S^{k^{n-1}1}$ by letting
$\psi(\bar{t})=[[\cdots[a_{0},a_{1,1},\dots,a_{1,n-1}],a_{2,1},\dots,a_{2,n-1}],\dots],a_{k,1},\dots,a_{k,n-1}].$
We claim that this map induces a well defined surjective $S_{k(n-1)+1}$-module homomorhism, $\psi:S^{k^{n-1}1}\to\rho_{n,k}/\tilde{\rho}_{n,k}$. To see this, first note that the column relations map to antisymmetry relations on brackets. The Garnir relations $g_{c}^{t}$ map to relations on combs that are the same as the relations obtained by setting the non-combs equal to 00 in the alternative Jacobi relation (2.1). Indeed, let $x_{1}$ in (2.1) be a comb and let $x_{2},\dots x_{n},y_{1},\dots,y_{n-1}$ be single letters, making the left hand side of (2.1) a comb. Then the summands of the right hand side of (2.1) are all combs except for the $i=1$ term, which is set equal to 00. Hence $\psi$ is well defined. Since, by Proposition 2.2, $\rho_{n,k}/\tilde{\rho}_{n,k}$ is spanned by combs, $\psi$ is surjective.
We claim that $\ker\psi=0$. Suppose not. Then since $S^{k^{n-1}1}$ is irreducible, $\ker\psi=S^{k^{n-1}1}$. It follows that $\rho_{n,k}=\tilde{\rho}_{n,k}$, which means that the non-combs span $\rho_{n,k}$. This implies that the non-combs of type $1^{k(n-2)+1}k$ span $\hat{\rho}_{n,k}$. Since the $b$-removing $S_{k(n-2)+1}$-module isomorphism $\varphi:\hat{\rho}_{n,k}\to\rho_{n-1,k}$ in the proof of (3.1) takes non-combs of type $1^{k(n-2)+1}k$ to non-combs of type $1^{k(n-2)+1}$, the non-combs span $\rho_{n-1,k}$. Continuing this argument leads to the conclusion that the non-combs span $\rho_{2,k}=\mathrm{Lie}_{k+1}$, which means that $\tilde{\rho}_{2,k}=\rho_{2,k}$. But by Pieri’s rule, all the irreducibles in $\tilde{\rho}_{2,k}$ have fewer than $k$ columns and thus the multiplicity of $S^{k^{1}1}$ in $\tilde{\rho}_{2,k}$ is 00, while by Part (1) of this theorem, the multiplicity of $S^{k^{1}1}$ in $\rho_{2,k}$ is $1$, which means that $\tilde{\rho}_{2,k}$ and $\rho_{2,k}$ cannot be equal. Thus the claim that $\ker\psi=0$ holds, from which we conclude that $\psi$ is indeed an isomorphism.
To prove (3) first note that when $k=1$ or $k=2$, there are no non-combs; so $\tilde{\rho}_{n,k}=0$. For the converse, let $\tilde{\rho}_{n,k}=0$. Then by part (2), $\rho_{n,k}=S^{k^{n-1}1}$. By Theorem 3.3, this implies $\rho_{2,k}=\mathrm{Lie}_{k+1}\cong S^{k^{1}1}$. We use the result of Kraskiewicz and Weyman discussed in the introduction to show that $k=1$ or $k=2$. Indeed, $\rho_{2,3}=S^{31}\oplus S^{21^{2}}$. Thus $k\neq 3$. If $k\geq 4$, consider the standard Young tableau of shape $(k-1)2$ whose first row is $1,2,4,5,\dots,k$ and whose second row is $13,k+$ 1. Since its major index is $2+k$, which is congruent to $11\bmod k+$ 1, we have that $S^{(k-1)^{1}2}$ is in $\rho_{2,k}$. Thus $k<4$. ∎
Note that Part (2) of the theorem generalizes Theorem 1.1 since there are no non-combs when $k=2$.
4. The non-comb submodule for $k=3$
Recall from Theorem 3.6 that $\tilde{\rho}_{n,3}$ denotes the submodule of $\rho_{n,3}$ spanned by the non-combs. In this section, we prove the following result, which by Theorem 3.6 is equivalent to Theorem 1.2.
Theorem 4.1.
For all $n\geq 2$,
$\tilde{\rho}_{n,3}\cong_{S_{3n-2}}S^{3^{n-2}21^{2}}.$
Let $\tilde{V}_{n,k}$ be the multilinear component of the vector space spanned by the $n$-ary non-combs on $[k(n-1)+1]$ subject only to the anticommuting relations, but not the generalized Jacobi relations. Clearly $\tilde{V}_{n,k}$ is invariant under the action of $S_{k(n-1)+1}$ on the multilinear component of the full vector space $V_{n,k}$ of bracketed permutations on $[k(n-1)+1]$ subject only to the anticommuting relations.
Now let $k=3$. For each $\alpha\in S_{3n-2}$, let $v_{\alpha}$ be the bracketed word (subject only to the anticommuting relations),
$[[\alpha(1),\dots,\alpha(n)],[\alpha(n+1),\dots,\alpha(2n)],\alpha(2n+1),\dots,\alpha(3n-2)].$
Clearly $\{v_{\alpha}:\alpha\in S_{3n-2}\}$ forms a set of generators for $\tilde{V}_{n,3}$ and $\sigma\in S_{3n-2}$ acts on the generators by $\sigma v_{\alpha}=v_{\sigma\alpha}$. Note that as $S_{3n-2}$-modules,
(4.1) $\tilde{V}_{n,3}\cong\left(\operatorname{sgn}_{2}[\operatorname{sgn}_{n}]\otimes\operatorname{sgn}_{n-2}\right)\uparrow_{S_{2n}\times S_{n-2}}^{S_{3n-2}},$
where $U[V]$ denotes the wreath product module obtained from the $S_{n}$-module $U$ and the $S_{m}$-module $V$. It is well known that the wreath product module $\operatorname{sgn}_{2}[\operatorname{sgn}_{n}]$ decomposes into irreducibles as follows
$\operatorname{sgn}_{2}[\operatorname{sgn}_{n}]\cong\bigoplus_{\begin{subarray}{c}i\in[n]\\ i\mbox{ \scriptsize{odd} }\end{subarray}}S^{2^{n-i}1^{2i}}.$
Hence
(4.2) $\tilde{V}_{n,3}\cong\bigoplus_{\begin{subarray}{c}i\in[n]\\ i\mbox{ \scriptsize{odd} }\end{subarray}}\left(S^{2^{n-i}1^{2i}}\otimes\operatorname{sgn}_{n-2}\right)\uparrow_{S_{2n}\times S_{n-2}}^{S_{3n-2}}.$
For $\alpha\in S_{3n-2}$, we form the relation
$R_{\alpha}:=2(n-2)v_{\alpha}-\sum_{\begin{subarray}{c}i\in[2n]\\ j\in[3n-2]\setminus[2n]\end{subarray}}v_{\alpha(i,j)}.$
It follows from antisymmetry that $R_{\alpha}$ is $2(n-2)$ times the relation
$v_{\alpha}-\sum_{i\in[n]}v_{\alpha(i,2n+1)},$
which by antisymmetry is the alternative generalized Jacobi relation (2.1). Consequently if $R$ is the submodule of $\tilde{V}_{n,3}$ spanned by $\{R_{\alpha}:\alpha\in S_{3n-2}\}$, then
(4.3) $\tilde{\rho}_{n,3}\subseteq\tilde{V}_{n,3}/R.$
We will show that the two modules are in fact equal to each other and isomorphic to the irreducible $S^{3^{n-2}21^{2}}$.
We begin by rewriting $R_{\alpha}$ in a more useful form:
(4.6) $R_{\alpha}$ $=2(n-2)v_{\alpha}-\sum_{\begin{subarray}{c}i\in[2n]\\ j\in[3n-2]\setminus[2n]\end{subarray}}v_{\alpha(i,j)}$
$=2(n-2)v_{\alpha}-\sum_{(i,j)\in S_{[3n-2]}}v_{\alpha(i,j)}+\sum_{(i,j)\in S_{[2n]}}v_{\alpha(i,j)}+\sum_{(i,j)\in S_{[3n-2]\setminus[2n]}}v_{\alpha(i,j)}$
$=\left(2(n-2)-\binom{n-2}{2}\right)v_{\alpha}-\sum_{(i,j)\in S_{[3n-2]}}v_{\alpha(i,j)}+\sum_{(i,j)\in S_{[2n]}}v_{\alpha(i,j)}$
(4.7) $=-\frac{(n-2)(n-7)}{2}v_{\alpha}-\sum_{(i,j)\in S_{[3n-2]}}(i,j)v_{\alpha}+\sum_{(i,j)\in S_{[2n]}}v_{\alpha(i,j)}$
with the third equality following from $v_{\alpha(i,j)}=-v_{\alpha}$ when $(i,j)\in S_{[3n-2]\setminus[2n]}$ which holds by antisymmetry of the bracket, and the fourth equality following from $\alpha(i,j)=(\alpha(i),\alpha(j))\alpha$.
We now consider the linear operators $I$ and $P$ on $\tilde{V}_{n,3}\cong\left(\operatorname{sgn}_{2}[\operatorname{sgn}_{n}]\otimes\operatorname{sgn}_{n-2}\right)\uparrow_{S_{2n}\times S_{n-2}}^{S_{3n-2}}$ defined on the elements $v$ of $\tilde{V}_{n,3}$ by
$I(v)=v\,\,\,\mbox{ and }\,\,\,P(v)=\sum_{(i,j)\in S_{[3n-2]}}(i,j)v.$
Let $Q$ be the linear operator defined on $\left(\operatorname{sgn}_{2}[\operatorname{sgn}_{n}]\otimes\operatorname{sgn}_{n-2}\right)\uparrow_{S_{2n}\times S_{n-2}}^{S_{3n-2}}$ by
$Q(u\otimes v)=\sum_{(a,b)\in S_{A}}(a,b)u\otimes v,$
where $u\in\operatorname{sgn}_{2}[\operatorname{sgn}_{n}]$ defined on the set $A\in\binom{[3n-2]}{2n}$ and $v\in\operatorname{sgn}_{n-2}$ defined on the complement $[3n-2]\setminus A$. The summands of (4.2) are easily seen to be invariant subspaces of $\tilde{V}_{n,3}$ under these operators.
For each odd $i$ in $[n]$, let $I_{i}$, $P_{i}$, and $Q_{i}$ be the restriction of $I$, $P$, and $Q$, respectively, to the summand $(S^{2^{n-i}1^{2i}}\otimes\operatorname{sgn}_{n-2})\uparrow_{S_{2n}\times S_{n-2}}^{S_{3n-2}}$ of $\tilde{V}_{n,3}$. It is well known that the operator $\sum_{(a,b)\in S_{n}}(a,b)$ acts on $S^{\lambda}$, for any partition $\lambda\vdash n$, by multiplication by the content $C(\lambda)=\sum_{(r,s)\in\lambda}c_{(r,s)}$, where $c_{(r,s)}=s-r$ for the cell in row $r$ and column $s$. Note that by Pieri’s rule, the irreducibles of $(S^{2^{n-i}1^{2i}}\otimes\operatorname{sgn}_{n-2})\uparrow_{S_{2n}\times S_{n-2}}^{S_{3n-2}}$ are multiplicity free and if $S^{\lambda}$ is such an irreducible, the skew shape $\lambda/(2^{n-i}1^{2i})$ is a vertical strip, that is, a skew shape in which no two cells are in the same row.
Lemma 4.2.
Let $i\in[n]$ be odd and let $S^{\lambda}$ be an irreducible of $(S^{2^{n-i}1^{2i}}\otimes\,\operatorname{sgn}_{n-2})\uparrow_{S_{2n}\times S_{n-2}}^{S_{3n-2}}$. Then
  1. (1)
    the operator $P_{i}$ acts on $S^{\lambda}$ as multiplication by the scalar $p_{i}(\lambda)=C(\lambda)$, and $p_{i}(\lambda)\leq C(2^{n-i}1^{2i})-\frac{(n-2)(n-7)}{2}$,
  2. (2)
    $p_{i}(\lambda)=C(2^{n-i}1^{2i})-\frac{(n-2)(n-7)}{2}$ if and only if $i=1$ and $\lambda=3^{n-2}21^{2}$,
  3. (3)
    the operator $Q_{i}$ acts on $S^{\lambda}$ as multiplication by the scalar $q_{i}=C(2^{n-i}1^{2i})$.
Proof of (1).
By the discussion prior to the statement of the lemma, we have that $P_{i}$ acts on $S^{\lambda}$ as multiplication by the scalar $C(\lambda)$. It follows from Pieri’s rule that $\lambda$ contains $2^{n-i}1^{2i}$ and that the cells of $\lambda$ that are not in $2^{n-i}1^{2i}$ form a vertical strip with at most $3$ columns. Clearly, we only need to show that the contribution to $C(\lambda)$ of the vertical strip is less than or equal to $-\frac{(n-2)(n-7)}{2}$.
Suppose the vertical strip consists of cells $(n+i+1,1),(n+i+2,1),\dots(n+i+j_{1},1)$ in column 1, cells $(n-i+1,2),(n-i+2,2),\dots(n-i+j_{2},2)$ in column 2, and cells $(1,3),(2,3),\dots,(j_{3},3)$ in column 3, where $j_{1}+j_{2}+j_{3}=n-2$. Then $n+i+1>n-i+j_{2}$ and $n-i+1>j_{3}$ since the vertical strip can’t have two cells in the same row. We rewrite this as
(4.8) $n+i\geq n-i+j_{2}\geq j_{2}+j_{3}.$
The contribution to $C(\lambda)$ of the vertical strip in column 1 is given by the left hand side of the inequality,
$j_{1}-(n+i+1)-(n+i+2)-\dots-(n+i+j_{1})\leq j_{1}-(j_{2}+j_{3}+1)-(j_{2}+j_{3}+2)-\dots-(j_{2}+j_{3}+j_{1}),$
which follows from (4.8). The contribution to $C(\lambda)$ of the vertical strip in column 2 is given by the left hand side of the inequality,
$2j_{2}-(n-i+1)-(n-i+2)-\dots-(n-i+j_{2})\leq 2j_{2}-(j_{3}+1)-(j_{3}+2)-\cdots-(j_{3}+j_{2}),$
which also follows from (4.8). The contribution to $C(\lambda)$ of the vertical strip in column 3 is given by
$3j_{3}-1-2-\dots-j_{3}.$
By summing these contributions we get that the contribution to $C(\lambda)$ of the vertical strip is less than or equal to
(4.9) $j_{1}+2j_{2}+3j_{3}-\!\!\sum_{k=1}^{j_{1}+j_{2}+j_{3}}k$ $\leq 3(j_{1}+j_{2}+j_{3})-\frac{(n-2)(n-1)}{2}$
$=3(n-2)-\frac{(n-2)(n-1)}{2}$
$=-\frac{(n-2)(n-7)}{2}.$
Proof of (2).
Note that the inequality of (4.9) is an equality if and only if $j_{1}=j_{2}=0$. In this case, the vertical strip is entirely in column 3 and the contribution to the content of $\lambda$ is precisely $-\frac{(n-2)(n-7)}{2}$. This means that the third column of $\lambda$ has length $n-2$, which implies that $n-i$, the length of the second column, is greater than or equal to $n-2$. Thus $i\leq 2$. Since $i$ is odd, $i=1$ and $\lambda=3^{n-2}21^{2}$. ∎
Proof of (3).
This follows from the fact that $Q_{i}$ is acting only on $S^{2^{n-i}1^{2i}}$ in $S^{2^{n-i}1^{2i}}\otimes\,\operatorname{sgn}_{n-2}$. ∎
We are now ready to prove Theorem 4.1 and thereby prove Theorem 1.2.
Proof of Theorem 4.1.
Let $\varphi=-\frac{(n-2)(n-7)}{2}I-P+Q.$ Then $\varphi$ is a linear operator on $\tilde{V}_{n,3}$. We claim that
(4.10) $\operatorname{Im}\varphi=R.$
Indeed, by (4.7), we have $\varphi(v_{\alpha})=R_{\alpha}$ for all $\alpha\in S_{3n-2}$. Since the $v_{\alpha}$ generate $\tilde{V}_{n,3}$ and the $R_{\alpha}$ generate $R$, the claim holds.
Next we find the kernel of $\varphi$. For each odd $i\in[n]$, let $\varphi_{i}=-\frac{(n-2)(n-7)}{2}I_{i}-P_{i}+Q_{i}.$ Then $\varphi_{i}$ is a linear operator on $(S^{2^{n-i}1^{2i}}\otimes\,\operatorname{sgn}_{n-2})\uparrow_{S_{2n}\times S_{n-2}}^{S_{3n-2}}$, and
$\ker(\varphi)=\bigoplus_{\mbox{ \scriptsize odd }i\in[n]}\ker(\varphi_{i}).$
It follows from parts (1) and (3) of Lemma 4.2 that if $S^{\lambda}$ is an irreducible of $(S^{2^{n-i}1^{2i}}\otimes\,\operatorname{sgn}_{n-2})\uparrow_{S_{2n}\times S_{n-2}}^{S_{3n-2}}$ then $\varphi_{i}$ acts on $S^{\lambda}$ as multiplication by the scalar
$-\frac{(n-2)(n-7)}{2}-p_{i}(\lambda)+q_{i}\geq 0.$
By part (2) we have that this scalar is 00 if and only if $i=1$ and $\lambda=3^{n-2}21^{2}$. It follows that $\ker\varphi=S^{3^{n-2}21^{2}.}$ We conclude from this and (4.10) that
$\tilde{V}_{n,3}/R\cong S^{3^{n-2}21^{2}}.$
Now by (4.3) we have that $\tilde{\rho}_{n,3}$ is either trivial or is isomorphic to the irreducible $S^{3^{n-2}21^{2}}$. The former case is impossible by part 3 of Theorem 3.6. Therefore $\tilde{\rho}_{n,3}\cong S^{3^{n-2}21^{2}}$, as desired. ∎
Acknowledgements
We thank Richard Stanley and Vic Reiner for helpful discussions.